OpenStudy (goformit100):

find the value of tan (-330) degrees

OpenStudy (goformit100):

@neer2890 @Koikkara @Kainui @iambatman

OpenStudy (goformit100):

Now let me start again

OpenStudy (koikkara):

\(\color{darkgreen}{~''Gof. ~write~ all~ your~ points~!!~then We~will~start"}\)

OpenStudy (goformit100):

tan (-330) = - tan(330) = - tan(90*3 + 60) = - cot(60) = -(-1/sqrt 3 ) = 1/sqrt 3 ......................... is this okay ?

OpenStudy (goformit100):

@neer2890 @Koikkara Check

OpenStudy (imstuck):

way too difficult! There's a way so much easier than that!!!!!

OpenStudy (goformit100):

@Kainui check my working

OpenStudy (kainui):

Just to be sure, what is your reasoning behind making this step: = - tan(90*3 + 60) = - cot(60)

OpenStudy (neer2890):

cot 60 \[\cot 60=\frac{ 1 }{ \sqrt{3} }\] i think you wrote-(cot 60) by mistaken instead of -(-cot 60)

OpenStudy (imstuck):

Use a reference angle....just sayin'! lalalalala.....

OpenStudy (anonymous):

Even easier is adding 360 to that angle to get 30 degrees and then doing tan(30)= sqrt(3)/3

OpenStudy (goformit100):

= - tan(90*3 + 60) this means by division method I have splitted it int as co-terminal angle

OpenStudy (imstuck):

Thank you doulikepiecauseidont!!!!!!!!!! That's the easy way!

OpenStudy (goformit100):

now = - cot(60) it is because since, tan (90 + 60) changes to it's co function. hence - cot 60

OpenStudy (goformit100):

Now after that value of cot 60 in forth quadrant will be - 1/sqrt of 3

OpenStudy (goformit100):

hence - ( - 1/sgrt 3) = 1/ sqrt 3 ......................... am I corrct guys ?

OpenStudy (goformit100):

*correct

OpenStudy (imstuck):

Yes you are correct...but it took way too much effort to get there. Good job, though! As long as you're correct that is all that matters.

OpenStudy (neer2890):

correct.. you can also use the method given by IMstuck. But what you did is the basic thing to remember.